# Constructing stable, high-order finite-difference operators on point clouds over complex geometries

### v1

@@ 5.1 Geometries, node distributions, and discretizations
Most of the studies in Sections 5.2, 5.3, and 5.4 share similar geometries, node distributions, and discretizations. These common elements are described in this section, to avoid repetition later.
5.1.1 Geometries: box-circle and annulus
Two geometries are adopted for the majority of the numerical studies. The first geometry is a unit square with a circle of radius $r=1/4$ excised from its center. This geometry was introduced in Figure 1(a) and is defined precisely below:
@@ 5.1.2 Node distributions
The node distributions that we use adopt depend on the geometry. For the box-circle geometry, we generate a uniform grid with $n_{x}=n_{y}$ nodes in each coordinate direction, and then randomly perturb the coordinates. We then remove all nodes that are inside the circle of radius 1/4. More precisely, consider the set of nodes $\hat{X}_{\text{box}}=\{\bm{x}_{i}\}_{i=1}^{n_{x}^{2}}$ where the coordinates of the $i$th node are given by
$x_{i}$ $=\big{(}j-{\textstyle\frac{1}{2}}\big{)}\Delta x+\xi_{i},$ $\forall j\in\{1,2,\ldots,n_{x}\}$
@@
$X_{\text{box}}=\{\bm{x}\in\hat{X}_{\text{box}}\;|\;\|\bm{x}-\bm{x}_{c}\|\geq{\textstyle\frac{1}{4}}\}.$
@@
$\rho_{i}$ $=g\left(\big{(}j-{\textstyle\frac{1}{2}}\big{)}\Delta\rho+\xi_{i}\right),$ $\forall j\in\{1,2,\ldots,n_{r}\}$
@@ 5.1.3 Norm-inequality tolerances
recall that $N$ depends on the particular node realization for $\Omega_{\text{box}}$. The annulus’ node distribution can be either quasi-uniform or clustered, so the following tolerance formula takes this into account:
@@ 5.2 Diagonal mass-matrix studies
This section focuses on characterizing the diagonal mass matrix and its construction. We begin by studying the distribution of entries in $\mathsf{M}$ before and after solving the linear inequality (18). Subsequently, we investigate the success rate for finding a positive-definite $\mathsf{M}$ for large sample of geometries and nodes. Finally, we verify the accuracy of $\mathsf{M}=\operatorname{\mathsf{diag}}(\bm{m})$ for numerical integration.
@@ 5.2.1 Quadrature weight distribution
The figures in the center column of Figure 4 depict the weights after enforcing the norm inequality. There are no white circles in Figures 4(b), 4(e), 4(h), and 4(k). The weights also show a clear pattern, increasing as we move from nodes near the inner radius to nodes near the outer radius.
@@ 5.2.2 Solution success rate
Recall that Theorem 5 does not guarantee that the norm inequality has a feasible solution, in general. It only implies that, for equidistributed $X_{N}$, there will be a feasible solution for sufficiently large $N$. Thus, the study presented in this section seeks to characterize what “sufficiently large” means in practice.
@@
The node generation process for each level-set $\phi_{k}(\bm{x})$ is similar to the one described in Section 5.1.1 for $\Omega_{\text{box}}$. We first create a quasi-uniform node distribution over the box $[-1,1]^{2}$ with $n_{x}=n_{y}$ nodes in each coordinate direction. We then add perturbations $\xi_{i}$ and $\eta_{i}$ to $x_{i}$ and $y_{i}$, respectively, where $\xi_{i},\eta_{i}\sim U[-\Delta x/4,\Delta x/4]$ and $\Delta x=2/n_{x}$. Finally, we remove nodes $i$ for which $\phi_{k}(\bm{x}_{i})<0$; if this produces fewer nodes than the minimum required by Algorithm 1, we add one to $n_{x}$ and repeat the process. For each geometry we consider four node resolutions corresponding to $n_{x}\in\{8,16,32,64\}$.
The tolerance in the norm inequality is based on the nominal node “volume” of $\Delta x^{2}=4/n_{x}^{2}$. As with the box-circle and annulus tolerances, we allow for quadrature weights that are one order of magnitude smaller than the nominal volume, so $\tau_{i}=4/(10n_{x}^{2})$.
@@
Figure 6 shows some conic geometries generated for the success-rate study. For $p=4$ on the coarsest node resolution ($n_{x}=8$), the norm inequality was successfully solved on geometries that are colored blue, and the inequality was not successfully solved for the red geometries. We emphasize that the coloring does not reflect the success for different $n_{x}$ or $p$. The figure is included primarily to illustrate the types of geometries that are generated and that there is no particular pattern in the failures for $p=4$ on the coarsest resolution.
@@
Figure 6: Example geometries used in the success-rate study; the white region indicates the domain $\Omega$. The $p=4$ norm was positive-definite on the coarsest mesh for the blue geometries, and it was negative-definite on the coarsest mesh for red geometries.
@@
Table 2 lists the percentage of geometries for which the norm-inequality problem was successfully solved. Each cell in the table is the success rate for a particular degree $p$ and resolution $n_{x}$ based on the 1000 sampled geometries. Two trends emerge from the data in the table.
@@
  • For a fixed node resolution, the norm inequality becomes increasingly infeasible as $p$ grows.
@@
  • For a fixed degree $p$, the probability of solving the norm inequality increases with the node resolution.
The second trend is consistent with the conclusions of Theorem 5.
@@
This study considered only quasi-uniform nodes, and the success rate is likely to decrease for strongly anisotropic node distributions. Nevertheless, regardless of the node distribution, this study and Theorem 5 suggest that adding nodes is a simple and effective remedy for an infeasible norm-inequality problem.
@@
$n_{x}$ 1 2 3 4
8 100.0% 100.0% 82.5% 8.1%
16 100.0% 100.0% 96.1% 77.7%
32 100.0% 100.0% 99.9% 97.9%
64 100.0% 100.0% 100.0% 100.0%
@@ 5.2.3 Norm quadrature accuracy
We assess the accuracy of the quadrature using both the box-circle and annulus domains. For the box-circle geometry, we consider five node resolutions with $n_{x}\in\{10,20,40,80,160\}$. For the annulus, we take $n_{\theta}=6n_{r}$ with $n_{r}\in\{6,12,24,48\}$. We consider two node distributions for the annulus: quasi-uniform ($\beta=0.1$) and radially clustered toward the inner circle ($\beta=4$). We generate 10 distinct node samples for each resolution, to avoid drawing conclusions based on outliers.
We exclude samples from the plots below if the norm inequality fails to be feasible. This was only necessary for samples from the annulus with the clustered node distribution: all 10 of the $p=4$ norm-inequality problems were infeasible on the coarsest grid, and 3 of the $p=3$ problems were infeasible on the coarsest grid.
@@
The numerical estimates for $I_{\text{box}}$ and $I_{\text{ann}}$ are computed using the formula
@@
The results of the quadrature accuracy study are presented in Figure 7. The absolute functional error is plotted versus the nominal mesh spacing for the box-circle (Fig. 7(a)), and the two node distributions for the annulus (Figs. 7(b) and 7(c)). The nominal mesh spacing, $h$, is based on the mean quadrature weight size:
@@
The dashed line in each convergence plot depicts the expected convergence rate of $\text{O}(h^{2p})$; this reference line is anchored to the smallest error in each data set.
Overall, the quadrature errors appear to converge at the expected rate of $2p$ or greater. The exceptions are the $p=4$ errors on the annulus with quasi-uniform nodes, and the $p=1$ errors on the annulus with clustered nodes.
@@ 5.3 Steady advection accuracy study
Having verified and characterized the diagonal mass matrix $\mathsf{M}$, we now turn our focus to SBP operators themselves. In this section, we use the steady advection equation to verify the solution accuracy of the SBP discretization (24).
@@
We adopt the same geometries and node distributions as the ones used in the quadrature-accuracy study (Section 5.2.3), except that we only consider the four coarser node resolutions for the box-circle geometry. For each sample, the $L^{2}$ error for the degree $p$ discretization is approximated using the corresponding SBP quadrature:
$L^{2}\;\text{error}\approx\sqrt{\big{(}\bm{u}-\bm{u}_{\text{exact}}\big{)}^{T}\mathsf{M}\big{(}\bm{u}-\bm{u}_{\text{exact}}\big{)}},$
@@
Figure 8 summarizes the results the SBP-accuracy study. The dashed reference lines in Figures 8(a), 8(b), and 8(c) show the design convergence rates of $p+1$. Comparing this reference line to the data at the finest and second-finest node resolutions, we conclude that the errors converge at the expected asymptotic rates. One exception is the $p=4$ error on the annulus with quasi-uniform nodes, which appears to be closer to a rate of 4 rather than 5.
@@
Figure 8: $L^{2}$ errors versus nominal mesh spacing for different node samples and geometries.

### v2

@@ 5.1 Geometries, node distributions, and discretizations
Most of the studies in Sections 5.25.6 share similar geometries, node distributions, and discretizations. These common elements are described in this section, to avoid repetition later.
5.1.1 Geometries: box-circle, annulus, and airfoil
Three geometries are adopted for the majority of the numerical studies. The first geometry is a unit square with a circle of radius $r=1/4$ excised from its center. This geometry was introduced in Figure 1(a) and is defined precisely below:
@@ 5.1.1 Geometries: box-circle, annulus, and airfoil
The third geometry is an airfoil shape that we include because it has a non-smooth feature. The airfoil is defined as
$\Omega_{\text{foil}}\equiv\big{\{}\bm{x}\in[0,1]\times[\textstyle{-\frac{1}{10}},\textstyle{\frac{1}{10}}]\;|\;x(x-1)^{2}-16y^{2}\geq 0\big{\}}.$
The rectangular domain $[0,1]\times[-1/10,1/10]$ in the definition of $\Omega_{\text{foil}}$ is chosen to produce a sharp trailing edge at $\bm{x}=[1,0]^{T}$ rather than a pair of intersecting curves. Figure 4 shows the airfoil geometry.
Figure 4: Airfoil geometry $\Omega_{\text{foil}}$ and an example node distribution used for the accuracy studies.
@@ 5.1.2 Node distributions
The node distributions that we use depend on the geometry. For the box-circle geometry, we generate a uniform grid with $n_{x}=n_{y}$ nodes in each coordinate direction, and then randomly perturb the coordinates. We then remove all nodes that are inside the circle of radius 1/4. More precisely, consider the set of nodes $\hat{X}_{\text{box}}=\{\bm{x}_{i}\}_{i=1}^{n_{x}^{2}}$ where the coordinates of the $i$th node are given by
$x_{i}$ $=\big{(}j-{\textstyle\frac{1}{2}}\big{)}\Delta x+\xi_{i},$ $\forall j\in\{1,2,\ldots,n_{x}\},$
@@
$X_{\text{box}}=\big{\{}\bm{x}\in\hat{X}_{\text{box}}\;|\;\|\bm{x}-\bm{x}_{c}\|\geq{\textstyle\frac{1}{4}}\big{\}}.$
@@
$\rho_{i}$ $=g\left(\big{(}j-{\textstyle\frac{1}{2}}\big{)}\Delta\rho+\xi_{i}\right),$ $\forall j\in\{1,2,\ldots,n_{r}\},$
@@
The nodes for the airfoil geometry are generated like the box-circle nodes, except there are five times more nodes in the $x$ direction than the $y$ direction. We begin by generating the set $\hat{X}_{\text{foil}}=\{\bm{x}_{i}\}_{i=1}^{5n_{y}^{2}}$ where the coordinates of $\bm{x}_{i}$ are given by
$x_{i}$ $=\big{(}j-{\textstyle\frac{1}{2}}\big{)}\Delta x+\xi_{i},$ $\forall j\in\{1,2,\ldots,5n_{y}\}$
$y_{i}$ $=-\frac{1}{10}+\big{(}k-{\textstyle\frac{1}{2}}\big{)}\Delta y+\eta_{i},$ $\forall k\in\{1,2,\ldots,n_{y}\},$
and $i=5kn_{y}+j$. The nominal node spacing is the same in the $x$ and $y$ directions: $\Delta x=\Delta y=1/(5n_{y})$. The random perturbations $\xi_{i}$ and $\eta_{i}$ are drawn from $U[-\Delta x/4,\Delta x/4]$, and the final node set is
$X_{\text{foil}}=\{\bm{x}\in\hat{X}_{\text{foil}}\;|\;x(x-1)^{2}-16y^{2}\geq 0\}.$
As with the box-circle, $N=|X_{\text{foil}}|$ depends on the sampled perturbations $\xi_{i}$ and $\eta_{i}$. Figure 4 shows a particular airfoil node distribution with $n_{y}=8$ and $N=220$ nodes total.
@@ 5.1.3 Norm-inequality tolerances
recall that $N$ depends on the particular node realization for $\Omega_{\text{box}}$.
The annulus’ node distribution can be either quasi-uniform or clustered, so the following tolerance formula takes this into account:
@@
Finally, we use the following tolerance for the airfoil geometry:
$\tau_{\text{foil},i}=\frac{1}{10}\Big{(}\frac{1}{5n_{y}}\Big{)}^{2},\qquad\forall\;i\{1,2,\ldots,N\}.$
Again, this represents one tenth of the average node “volume,” since the nominal node spacing in the $x$ and $y$ directions is $1/(5n_{y})$ for the airfoil geometry.
Remark 7.
Setting $\tau_{i}$ to one tenth the average node “volume” is informed by the results of the success-rate study in Section 5.2.2. However, this heuristic will likely fail for some node distributions, particularly those that are highly nonuniform and anisotropic, so developing a general algorithm to set $\tau_{i}$ will be the focus of future work.
@@ 5.2 Diagonal mass-matrix studies
This section focuses on characterizing the diagonal mass matrix and its construction. We begin by studying the distribution of entries in $\mathsf{M}$ before and after solving the linear inequality (18). Subsequently, we investigate the success rate for finding a positive-definite $\mathsf{M}$ for a large sample of geometries and nodes. Finally, we verify the accuracy of $\mathsf{M}=\operatorname{\mathsf{diag}}(\bm{m})$ for numerical integration.
@@ 5.2.1 Quadrature weight distribution
Remark 8.
The quasi-uniform nodes are not equidistributed — despite being quasi-uniform in $(r,\theta)$ space — because the nodes become more sparsely distributed as we move away from the center of the annulus. Consequently, we should expect the quadrature weights to vary in the radial direction for both the $\beta=0.1$ and the $\beta=4$ cases.
@@
The figures in the center column of Figure 5 depict the weights after enforcing the norm inequality. There are no white circles in Figures 5(b), 5(e), 5(h), and 5(k). The weights also show a clear pattern, increasing as we move from nodes near the inner radius to nodes near the outer radius.
@@ 5.2.2 Solution success rate
Recall that Theorem 5 does not guarantee that the norm inequality has a feasible solution, in general. It only implies that, for equidistributed $X_{N}$, there will be a feasible solution for sufficiently large $N$. Thus, the study presented in this section seeks to characterize what “sufficiently large” means in practice. The study also investigates the role that the tolerances $\tau_{i}$ play in determining feasibility.
@@
The node generation process for each level-set $\phi_{k}(\bm{x})$ is similar to the one described in Section 5.1.1 for $\Omega_{\text{box}}$ and $\Omega_{\text{foil}}$. We first create a quasi-uniform node distribution over the box $[-1,1]^{2}$ with $n_{x}=n_{y}$ nodes in each coordinate direction. We then add perturbations $\xi_{i}$ and $\eta_{i}$ to $x_{i}$ and $y_{i}$, respectively, where $\xi_{i},\eta_{i}\sim U[-\Delta x/4,\Delta x/4]$ and $\Delta x=2/n_{x}$. Finally, we remove nodes $i$ for which $\phi_{k}(\bm{x}_{i})<0$; if this produces fewer nodes than the minimum required by Algorithm 1, we add one to $n_{x}$ and repeat the process. For each geometry we consider four node resolutions corresponding to $n_{x}\in\{8,16,32,64\}$.
As mentioned above, we are also interested in how the tolerances $\tau_{i},i=1,2,\dots,N$ impact feasibility of the norm inequality. Intuitively, $\tau_{i}$ should be smaller than the average node “volume” $\Delta x^{2}=4/n_{x}^{2}$; however, other than this upper bound on the $\tau_{i}$, the relationship between the tolerances and feasibility is not immediately obvious. To help elucidate this relationship, we consider three values for $\tau_{i}$:
$\tau^{\text{large}}=\frac{1}{(n_{x})^{2}},\qquad\tau^{\text{small}}=\frac{1}{(10n_{x})^{2}},\qquad\text{and}\qquad\tau^{\text{tiny}}=\frac{1}{(100n_{x})^{2}}.$
The large tolerance is one quarter of the average node “volume,” while the small and tiny tolerances are 1/400 and 1/40000 of the average, respectively.
@@
Figure 7 shows some conic geometries generated for the success-rate study. For $p=4$ on the coarsest node resolution ($n_{x}=8$) and $\tau_{i}=\tau_{i}^{\text{small}}$, the norm inequality was successfully solved on geometries that are colored blue, and the inequality was not successfully solved for the red geometries. We emphasize that the coloring does not reflect the success for different $n_{x}$ or $p$. The figure is included primarily to illustrate the types of geometries that are generated and that there is no particular pattern in the failures for $p=4$ on the coarsest resolution.
@@
Figure 7: Example geometries used in the success-rate study; the white region indicates the domain $\Omega$. Using $\tau_{i}=\tau^{\text{small}}$, the $p=4$ norm was positive-definite on the coarsest mesh for the blue geometries, and it was negative-definite on the coarsest mesh for red geometries.
@@
Table 2 lists the percentage of geometries for which the norm-inequality problem was successfully solved. Each cell in the table is the success rate for a particular tolerance $\tau_{i}$, degree $p$, and resolution $n_{x}$ based on the 1000 sampled geometries. Three trends emerge from the data in the table.
  • For a fixed degree $p$ and node resolution $n_{x}$, decreasing the tolerance helps improve feasibility, up to a point. In particular, we see a dramatic improvement in success rate in going from $\tau^{\text{large}}$ to $\tau^{\text{small}}$, but limited improvement in success rate when decreasing the tolerance from $\tau^{\text{small}}$ to $\tau^{\text{tiny}}$.
@@
  • For a fixed tolerance and node resolution, the norm inequality becomes increasingly infeasible as $p$ grows.
@@
  • For a fixed tolerance and degree $p$, the probability of solving the norm inequality increases with the node resolution.
The third trend is consistent with the conclusions of Theorem 5.
@@
This study considered only quasi-uniform nodes, and the success rate is likely to decrease for strongly anisotropic node distributions. Nevertheless, regardless of the node distribution, this study and Theorem 5 suggest that adding nodes is a simple and effective remedy for an infeasible norm-inequality problem. In addition, the study suggests that there exists a threshold for $\tau_{i}$ below which feasibility is relatively insensitive.
@@
$\tau_{i}$ $n_{x}$ 1 2 3 4
$\tau_{i}^{\text{large}}=1/(n_{x})^{2}$ 8 92.1% 89.1% 55.9% 2.0%
16 99.9 % 99.9% 89.9% 71.9%
32 100.0% 100.0% 99.5 % 96.2%
64 100.0% 100.0% 100.0% 99.9%
$\tau_{i}^{\text{small}}=1/(10n_{x})^{2}$ 8 100.0% 100.0% 79.7% 9.6%
16 100.0% 100.0% 92.9% 75.5%
32 100.0% 100.0% 99.6% 97.2%
64 100.0% 100.0% 100.0% 99.9%
$\tau_{i}^{\text{tiny}}=1/(100n_{x})^{2}$ 8 100.0% 100.0% 80.0% 9.7%
16 100.0% 100.0% 92.9% 75.8%
32 100.0% 100.0% 99.6% 97.2%
64 100.0% 100.0% 100.0% 99.9%
@@ 5.2.3 Norm quadrature accuracy
We assess the accuracy of the quadrature using the box-circle, annulus, and airfoil domains. For the box-circle geometry, we consider five node resolutions with $n_{x}\in\{10,20,40,80,160\}$. For the annulus, we take $n_{\theta}=6n_{r}$ with $n_{r}\in\{6,12,24,48\}$. We consider two node distributions for the annulus: quasi-uniform ($\beta=0.1$) and radially clustered toward the inner circle ($\beta=4$). The airfoil geometry uses node resolutions defined by $n_{y}\in\{4,8,16,32,64\}$; recall that $n_{x}=5n_{y}$ for the airfoil. We generate 10 distinct node samples for each geometry and resolution, to avoid drawing conclusions based on outliers.
We exclude samples from the plots below if the norm inequality fails to be feasible. This was necessary for some samples from the annulus with the clustered node distribution and some samples from the airfoil geometry. Specifically, three samples from the annulus and two samples from the airfoil geometry failed for $p=3$ on the coarsest distributions, and all ten samples for $p=4$ failed on the coarsest distributions for both problems.
@@
and the integral used on the airfoil domain is
[…]
where $\operatorname{\mathsf{erfi}}(x)$ is the imaginary error function, which we evaluate using the SpecialFunctions.jl package. The numerical estimates for $I_{\text{box}}$, $I_{\text{ann}}$, and $I_{\text{foil}}$ are computed using the formula
@@
The results of the quadrature accuracy study are presented in Figure 8. The absolute functional error is plotted versus the nominal mesh spacing for the box-circle (Fig. 8(a)), the two node distributions for the annulus (Figs. 8(b) and 8(c)), and the airfoil (Fig. 8(d)). The nominal mesh spacing, $h$, is based on the mean quadrature weight size:
@@
The solid lines in the convergence plots indicate the expected asymptotic behavior for the functional errors, i.e., $\text{O}(h^{2p})$. These reference lines are colored based on the corresponding $p$ and are anchored to the smallest error in each data set.
Overall, the quadrature errors appear to converge at the expected rate of $2p$ or greater. The exceptions are the $p=4$ errors on the annulus with quasi-uniform nodes; the $p=1$ errors on the annulus with clustered nodes; and the $p=3$ and $p=4$ errors on the airfoil, where round-off errors are impacting the results on the finest node distributions.
@@ 5.3 Steady advection accuracy study
Having verified and characterized the diagonal mass matrix $\mathsf{M}$, we now turn our focus to the SBP operators themselves. In this section, we use the steady advection equation to verify the solution accuracy of the SBP discretization (24).
@@
We adopt the same geometries and node distributions as the ones used in the quadrature-accuracy study (Section 5.2.3), except that we only consider the four coarser node resolutions for the box-circle and airfoil geometries. For each sample, the $L^{2}$ error for the degree $p$ discretization is approximated using the corresponding SBP quadrature:
$L^{2}\;\text{error}\approx\sqrt{\big{(}\bm{u}-\bm{u}_{\text{exact}}\big{)}^{T}\mathsf{M}\big{(}\bm{u}-\bm{u}_{\text{exact}}\big{)}},$ (26)
@@
Figure 9(d) summarizes the results of the SBP-accuracy study. The reference lines in Figures 9(a), 9(b), 9(c), and 9(d) show the design convergence rates of $p+1$, and the lines are anchored to the smallest error from the corresponding degree $p$ data set. Comparing these reference lines to the data at the finest and second-finest node resolutions, we conclude that the errors converge at the expected asymptotic rates for most of the geometries and distributions, with the following exceptions: the $p=4$ error on the annulus with quasi-uniform nodes appears to be closer to a rate of four rather than five; the $p=1$ airfoil discretization appears to be first-order only, and; the $p=4$ airfoil discretization is between third and fourth order.
@@
(d) airfoil
Figure 9: $L^{2}$ errors versus nominal mesh spacing for different node samples and geometries from the steady advection accuracy study
5.4 Efficiency study
Point-cloud SBP operators are less sparse than classical (tensor-product) SBP finite-difference operators, which has implications for the relative efficiency of the point-cloud discretizations. To help quantify their efficiency, this section compares the performance of point-cloud discretizations with that of tensor-product SBP discretizations applied to the linear advection problem.
We adopt the annulus domain for the efficiency study, because it simplifies mesh generation for the classical SBP operators. The tensor-product discretization uses a conforming grid with node coordinates given by $\bm{x}_{i}=[r_{i}\cos(\theta_{i}),r_{i}\sin(\theta_{i})]^{T}$, where
$r_{i}$ $=\frac{1}{2(n_{r}-1)}\big{(}n_{r}-1+j-1\big{)},$ $\qquad\forall\,j$$\in\{1,2,\ldots,n_{r}\},$
$\theta_{i}$ $=\frac{2\pi(k-1)}{n_{\theta}},$ $\qquad\forall\,k$$\in\{1,2,\ldots,n_{\theta}\},$
and $i=kn_{r}+j$. For the radial direction we take $n_{r}\in\{11,21,41,81,161\}$ and for the angular direction $n_{\theta}=6(n_{r}-1)+1$.
The tensor-product discretization is standard. In the radial direction we use the second- ($p=1$), third- ($p=2$), fourth- ($p=3$), and fifth- ($p=4$) order classical SBP operators from Mattsson and Nordström [34] as implemented in the SummationByPartsOperators.jl package [35]. Grids for which $n_{r}$ is too small for the given SBP operators are excluded from the study. In the angular direction we use centered, periodic finite-difference operators of order $2p$; these periodic operators have the same finite-difference coefficients as the interior stencils of the radial SBP operators. We use upwind numerical flux functions along the boundaries $r=1/2$ and $r=1$ to impose the boundary conditions. Finally, we compute the terms in the mapping Jacobian using the appropriate finite-difference operator, following the typical tensor-product discretization in $(r,\theta)$ space; see, for example, [36].
The nodes for the point-cloud SBP operators are generated as described in Section 5.1.2, and they use the quasi-uniform ($\beta=0.1$) distribution with $n_{\theta}=6n_{r}$ and $n_{r}\in\{6,12,24,48\}$. We show results for only one node-set realization from Section 5.3 to avoid cluttering the figures, but the results for the other nine realizations are similar.
For each discretization, we solve the linear advection problem based on the manufactured solution given in Section 5.3 and compute the $L^{2}$ error using Equation (26). Figure 10 shows the $L^{2}$ errors plotted versus the number of nodes, $N$, as well as the number of non-zeros ($nnz$) in the system matrix. We use $nnz$ as a proxy for computational cost, since the point-cloud implementation has not been optimized.
The first (left) plot in Figure 10 shows that, for this problem, the $L^{2}$ errors from the two discretizations overlap with one another when plotted versus number of nodes. However, the second (right) plot shows that the classical SBP discretization is significantly more efficient with respect to $nnz$. Indeed, for the same $L^{2}$ error and degree $p$, the classical SBP discretization requires between four ($p=1$) and ten ($p=4$) times fewer non-zeros in the system matrix.
Figure 10: $L^{2}$ error versus number of nodes (left figure) and $L^{2}$ error versus the number of non-zeros in the advection system matrix (right figure). Results are shown for both classical and point-cloud SBP discretizations of the advection problem on the annulus domain.